### v1

# LS-category and sequential topological complexity of symmetric products

Abstract.
The $n$-th symmetric product of a topological space $X$ is the orbit space of the natural action of the symmetric group $S_{n}$ on the product space $X^{n}$. In this paper, we compute the sequential topological complexities of the symmetric products of closed orientable surfaces, thereby verifying the rationality conjecture of Farber and Oprea for these spaces. Additionally, we determine the Lusternik–Schnirelmann category of the symmetric products of closed non-orientable surfaces. More generally, we provide lower bounds to the LS-category and the sequential topological complexities of the symmetric products of finite CW complexes $X$ in terms of the cohomology of $X$ and its products. On the way, we also obtain new lower bounds to the sequential distributional complexities of continuous maps and study the homotopy groups of the symmetric products of closed surfaces.
Key words and phrases:
Symmetric product, sequential topological complexity, LS-category, finite CW complex, homotopy group, sequential distributional complexity.
2010 Mathematics Subject Classification:
Primary 55M30, 55S15, Secondary 57M20, 55N10.
1. Introduction
The topological complexity of a topological space [Far][Ru] is a homotopy invariant that measures the complexity of moving points continuously within the space. More precisely, given a topological space $Y$ and integer $m\geq 2$, the $m$-th sequential topological complexity of $Y$, denoted $\operatorname{\mathsf{TC}}_{m}(Y)$, measures the complexity of constructing a continuous motion planning algorithm (for mechanical systems having configuration space $Y$) that takes as input ordered $m$-tuples of positions in $Y$ and produces as output a uniformly timed motion of the system in $Y$ through those $m$ positions attained in that order. If $Y$ is the $n$-th symmetric product of a topological space $X$, denoted $SP^{n}(X)$, then positions in $Y$ become unordered $n$-tuples of positions in $X$.
Though the notion of sequential topological complexity of spaces was initially motivated by topological robotics, it has been studied as a homotopy invariant for several classes of spaces, see, for example, [Far][Ru][BGRT][LS][FO][HL], and the references therein. Furthermore, it is closely related to another homotopy invariant of spaces, called the Lusternik–Schnirelmann category ($\operatorname{\mathsf{cat}}$), which is much older, [BG][Ja][CLOT]. These homotopy invariants are well-known to be special cases of Schwarz’s general notion of the sectional category of fibrations, [Sch]. We refer the reader to [CLOT][Far][Ru] for details.
In general, precisely determining the LS-category of a space is difficult, and determining (or even estimating) its (sequential) topological complexity can be even more difficult, see, for example, [FTY][Dr2][CV].
In this paper, we study the problem of estimating the sequential topological complexities and the Lusternik–Schnirelmann category of symmetric products of topological spaces, specifically finite CW complexes. The spaces of our particular interest are the symmetric products of closed $2$-manifolds (or surfaces), which are closed even-dimensional manifolds.
Our first main result is the exact computation of the sequential topological complexities of the symmetric products of closed orientable surfaces $M_{g}$ of genus $g\geq 0$. The precise statement of our main result is as follows.
Theorem A.
For any integers $n,m\geq 2$ and $g\geq 0$, we have for the $m$-th sequential topological complexity of the $n$-th symmetric product of $M_{g}$ that
$\operatorname{\mathsf{TC}}_{m}\left(SP^{n}(M_{g})\right)=m\operatorname{\mathsf{cat}}\left(SP^{n}(M_{g})\right)=\begin{cases}2mn&\text{ if }n\leq g\\ mn+mg&\text{ if }n>g.\end{cases}$
To prove Theorem A, we use the description of the rational cohomology ring of $SP^{n}(M_{g})$ due to Macdonald [Mac], and we show that for each $m\geq 2$, the cohomological lower bound to $\operatorname{\mathsf{TC}}_{m}(SP^{n}(M_{g}))$ coincides with the LS-categorical upper bound, thereby determining $\operatorname{\mathsf{TC}}_{m}(SP^{n}(M_{g}))$. See Section 2.B for some preliminary facts on the bounds to sequential topological complexities.
The sequential topological complexities of orientable surfaces were computed in [Ru][BGRT] and [LS], and [GGGHMR]. The topological complexity ($\operatorname{\mathsf{TC}}_{2}$) of symmetric products of orientable surfaces was computed very recently in [DDJ]. Theorem A extends the work of [DDJ] by completing the computation of all the sequential topological complexities of all symmetric products of orientable surfaces. Furthermore, it verifies the rationality conjecture [FO] for the symmetric products of orientable surfaces.
The rationality conjecture in its original form can be stated as follows.
1.1 Conjecture (Farber–Oprea).
For a finite CW complex $X$, the formal power series
[…]
of the $\operatorname{\mathsf{TC}}$-generating function of $X$ represents a rational function of the form
$\frac{P_{X}(t)}{(1-t)^{2}},$
where $P_{X}(t)$ is an integral polynomial with $P_{X}(1)=\operatorname{\mathsf{cat}}(X)$.
This conjecture is known for only a few classes of finite CW complexes and closed manifolds (see Section 5.B), and its second part, which states that $P_{X}(1)=\operatorname{\mathsf{cat}}(X)$, has recently been disproven in [FKS] using a finite $11$-dimensional CW complex. However, the original conjecture is still open for closed manifolds.
Using Theorem A, we show that Conjecture 1.1 is verified for the symmetric products of orientable surfaces. This extends the class of closed manifolds for which the rationality conjecture holds.
Corollary A.1.
For any integers $n\geq 2$ and $g\geq 0$, the formal power series of the $\operatorname{\mathsf{TC}}$-generating function of the closed orientable $2n$-manifold $SP^{n}(M_{g})$ represents the rational function
$\frac{(2t-t^{2})\operatorname{\mathsf{cat}}(SP^{n}(M_{g}))}{(1-t)^{2}}.$
Another consequence of Theorem A is the following.
Corollary A.2.
For any integers $m,n\geq 1$ and $g\geq 0$, we have that
$\operatorname{\mathsf{cat}}((SP^{n}(M_{g}))^{m})=m\operatorname{\mathsf{cat}}(SP^{n}(M_{g})).$
In [LS], Lupton and Scherer proved that $\operatorname{\mathsf{TC}}_{m}(X)=\operatorname{\mathsf{cat}}(X^{m-1})$ if $X$ is an $H$-space, and they asked for examples of spaces $X$ for which $\operatorname{\mathsf{TC}}_{m}(X)=\operatorname{\mathsf{cat}}(X^{m})$ for each $m\geq 2$. Due to the above results, the manifolds $SP^{n}(M_{g})$ serve as examples of this kind for all $n$ and $g$, except for the case $n=g=1$ when the manifold is an $H$-space, namely the $2$-torus.
After studying the symmetric products of orientable surfaces, we turn towards the symmetric products of closed non-orientable surfaces $N_{g}$ of genus $g\geq 1$, which, compared to their orientable counterparts, have been studied significantly less in the literature. Our next main result is the exact computation of the LS-category of the symmetric products of non-orientable surfaces. The precise statement of our result is as follows.
Theorem B.
For any integers $n,g\geq 1$, we have for the Lusternik–Schnirelmann category of the $n$-th symmetric product of $N_{g}$ that
$\operatorname{\mathsf{cat}}(SP^{n}(N_{g}))=2n.$
Theorem B is proved by studying the ${\mathbb{Z}}_{2}$-cohomology ring of $SP^{n}(N_{g})$, which was described by Kallel and Salvatore in [KS]. Basically, we show that the ${\mathbb{Z}}_{2}$-cup-length of $SP^{n}(N_{g})$ coincides with the dimension of $SP^{n}(N_{g})$.
The LS-category of the symmetric products of orientable surfaces, as mentioned in Theorem A, can be determined using some results of [DDJ] and [Dr3] (see Section 5.A). Hence, Theorem B completes the computation of the LS-category of all symmetric products of all closed surfaces, thereby giving a complete list of closed surfaces $S$ for which $SP^{n}(S)$ is essential, in the sense of Gromov, [Gr].
Corollary B.1.
Given a closed surface $S$ and integer $n\geq 1$, the $n$-th symmetric product $SP^{n}(S)$ of $S$ is an essential $2n$-manifold if and only if either $S$ is an orientable surface of genus $\geq n$ or $S$ is a non-orientable surface of genus $\geq 1$.
In this paper, more generally, we study lower bounds to $\operatorname{\mathsf{TC}}_{m}(SP^{n}(X))$ and $\operatorname{\mathsf{cat}}(SP^{n}(X))$ for finite CW complexes $X$.
While the LS-category of products of a space $Y$ typically gives tight bounds to $\operatorname{\mathsf{TC}}_{m}(Y)$, the most useful lower bounds to $\operatorname{\mathsf{TC}}_{m}(Y)$ come from the cohomology of $Y$ and $Y^{m}$. Indeed, for any ring $R$, the $m$-th $R$-zero-divisor cup-length of $Y$, denoted $zc\ell^{m}_{R}(Y)$, is a lower bound to $\operatorname{\mathsf{TC}}_{m}(Y)$. We recall that $zc\ell_{R}^{m}(Y)$ is the cup-length of the kernel of the homomorphism induced in $R$-cohomology by the diagonal map $\Delta_{m}:Y\to Y^{m}$. If $Y$ is a symmetric product $SP^{n}(X)$ of a space $X$, which is the regime of our interest, then its cohomology ring can be complicated. In particular, computing the $m$-th $R$-zero-divisor cup-length of $SP^{n}(X)$ can be quite difficult. Therefore, we want to provide a useful lower bound to $\operatorname{\mathsf{TC}}_{m}(SP^{n}(X))$ in terms of the cohomology of $X$ and $X^{m}$.
It turns out that for any finite CW complex $X$, the rational cup-length and the $m$-th rational zero-divisor cup-length of $X$ give tight bounds to $\operatorname{\mathsf{cat}}(SP^{n}(X))$ and $\operatorname{\mathsf{TC}}_{m}(SP^{n}(X))$, respectively. In fact, they offer sharp lower bounds to some other homotopy invariants of interest that bound $\operatorname{\mathsf{cat}}(SP^{n}(X))$ and $\operatorname{\mathsf{TC}}_{m}(SP^{n}(X))$ from below. This observation is encapsulated in the following statement.
Theorem C.
For any finite CW complex $X$, ring $R$, and integer $n\geq 1$, we have that $c\ell_{R}(X)\leq c\ell_{R}(SP^{n}(X))$ and $c\ell_{{\mathbb{Q}}}(X)=c\ell_{{\mathbb{Q}}}(\textup{Im}(\delta_{n}^{*}))$. Moreover, for any integer $m\geq 2$, we have that $zc\ell^{m}_{{\mathbb{Q}}}(\textup{Im}((\delta_{n}^{m})^{*}))=zc\ell^{m}_{{\mathbb{Q}}}(X)\leq zc\ell_{{\mathbb{Q}}}^{m}(SP^{n}(X))$.
Here, $\delta_{n}:X\to SP^{n}(X)$ is the diagonal embedding defined as $\delta_{n}(x)=[x,\ldots,x]$, $\delta_{n}^{m}:X^{m}\to(SP^{n}(X))^{m}$ is the product map defined as
$\delta_{n}^{m}(x_{1},\ldots,x_{m})=\left(\delta_{n}(x_{1}),\ldots,\delta_{n}(x_{m})\right),$
and given the diagonal map $\Delta_{m}:X\to X^{m}$, the quantity $zc\ell^{m}_{{\mathbb{Q}}}(\textup{Im}(\left(\delta_{n}^{m}\right)^{*}))$ is defined as
$zc\ell^{m}_{{\mathbb{Q}}}(\textup{Im}(\left(\delta_{n}^{m}\right)^{*})):=c\ell_{{\mathbb{Q}}}\left(\textup{Im}(\left(\delta_{n}^{m}\right)^{*})\cap\text{Ker}\left(\Delta_{m}^{*}\right)\right).$
There is another lower bound to the $m$-th sequential topological complexity of a space, namely its $m$-th sequential distributional complexity $(\operatorname{\mathsf{dTC}}_{m})$, which is a new homotopy invariant of probabilistic flavor that also has some very interesting interpretations in terms of motion planning, see [DJ][KW][J1].
While studying lower bounds to sequential distributional complexities of CW complexes, symmetric products of these spaces and their cohomology rings came into the picture. More precisely, some general cohomological lower bounds to $\operatorname{\mathsf{dTC}}_{m}(X)$ were found in [J1] (see [DJ] for $m=2$) in terms of the cohomology of $SP^{n}(X)$ and $SP^{n}(X^{m})$ using the diagonal embeddings $X^{m}\to SP^{n}(X^{m})$. So, it becomes very tempting to explore relationships between lower bounds to the $m$-th sequential distributional complexity of spaces and the $m$-th sequential topological complexity of their diagonal embeddings into symmetric products for each $m\geq 2$. In that direction, we show that these two homotopy invariants share a common cohomological lower bound, namely the $m$-th rational zero-divisor cup-length of the space. We do this by finding sharp lower bounds to sequential distributional complexities of continuous maps. The precise statement of our general result in Alexander–Spanier cohomology is as follows.
Theorem D.
Let $f:X\to Y$ be a continuous map, $m\geq 2$, and $R$ be a ring with unity. For $1\leq i\leq n$ and some integers $k_{i}\geq 1$, let $\alpha_{i}^{*}\in H^{k_{i}}((SP^{n!}(Y))^{m};R)$ be $m$-th $R$-zero-divisors of $SP^{n!}(Y)$. Let $\alpha_{i}\in H^{k_{i}}(X^{m};R)$ be their images under $(\delta_{n!}^{m}\circ f^{m})^{*}:H^{*}((SP^{n!}(Y))^{m};R)\to H^{*}(X^{m};R)$. If $\alpha_{1}\smile\cdots\smile\alpha_{n}\neq 0$, then we have for the $m$-th sequential distributional complexity of the map $f$ that
$\operatorname{\mathsf{dTC}}_{m}(f)\geq n.$
While a cohomological lower bound to $\operatorname{\mathsf{dTC}}_{m}(f)$ can be readily obtained in terms of the cohomology of $SP^{k}(Y^{m})$ by generalizing the proofs of [J1, Section 4] from spaces to maps, the relevance of the lower bound obtained in Theorem D is that in general, the cohomology of $(SP^{k}(Y))^{m}$ is simpler (i)(i)(i) To see this, take $Y$ to be a circle or a closed surface and use Kunneth formulae — see [Mac] and Section 5.A in the case of orientable surfaces and [KS] and Section 6.B otherwise. than that of $SP^{k}(Y^{m})$ when $m,k\geq 2$. Hence, Theorem D makes it easier to bound $\operatorname{\mathsf{dTC}}_{m}(f)$ from below. Furthermore, it helps obtain the following convenient lower bound to $\operatorname{\mathsf{dTC}}_{m}(X)$.
Corollary D.1.
For any finite CW complex $X$ and integer $m\geq 2$, we have for the $m$-th sequential distributional complexity of $X$ that $zc\ell^{m}_{{\mathbb{Q}}}(X)\leq\operatorname{\mathsf{dTC}}_{m}(X)$.
Corollary D.1 subsumes the computations of $\operatorname{\mathsf{dTC}}_{m}$ for orientable surfaces and products of spheres done in [J1, Section 7], which are rather long and complicated. It also implies the following for the symmetric products of closed orientable surfaces.
Corollary D.2.
For any integers $m\geq 2$, $n\geq 1$, and $g\geq 0$, we have the equalities $\operatorname{\mathsf{dTC}}_{m}(SP^{n}(M_{g}))=\operatorname{\mathsf{TC}}_{m}(SP^{n}(M_{g}))$ and $\operatorname{\mathsf{dcat}}((SP^{n}(M_{g}))^{m})=\operatorname{\mathsf{cat}}((SP^{n}(M_{g}))^{m}).$
Some other results in this paper on the homotopy of the symmetric products of closed surfaces could be of interest on their own.
For a surface $S$, the fundamental group of $SP^{n}(S)$ for each $n\geq 2$ is the first integral homology group of $S$, see, for example, [KT]. The second homotopy group of $SP^{n}(M_{g})$ was studied in [BR], and its computation for $n\geq 3$ was recovered in [DDJ] recently using a cellular complex construction from [KS]. Here, using the Dold–Thom homotopy splitting [DT] of the infinite symmetric product of some finite $2$-dimensional CW complexes and the work of [KS], we show that the higher homotopy groups of the finite symmetric products of these spaces vanish up to a certain stage. This gives the following result for the symmetric products of surfaces, which also recovers the computation of $\pi_{2}(SP^{n}(M_{g}))$ for $n\geq 3$.
Theorem E.
Given a closed surface $S$ and integer $n\geq 3$, the $i$-th homotopy group of the $n$-th symmetric product of $S$ vanishes for each $3\leq i<n$. Moreover, for the second homotopy group, we have for each $n\geq 3$ and genus $g$ that
$\pi_{2}(SP^{n}(M_{g}))={\mathbb{Z}}\hskip 8.53581pt\text{and}\hskip 8.53581pt\pi_{2}(SP^{n}(N_{g}))=0.$
In particular, while $M_{g-1}$ is the orientable double cover of $N_{g}$ for each $g\geq 1$, a finite cover of $SP^{n}(M_{h})$ cannot be a finite cover of $SP^{n}(N_{g})$ for any choice of integers $n\geq 3$, $g\geq 1$, and $h\geq 0$.
In [DDJ], it was shown that the universal cover of the orientable manifold $SP^{n}(M_{g})$ admits a spin structure if and only if $n-g$ is odd. In contrast, as a result of Theorem E, we conclude the following in the non-orientable setting.
Corollary E.1.
For any integers $n\geq 3$ and $g\geq 1$, the universal cover $\widetilde{SP^{n}(N_{g})}$ of the non-orientable manifold $SP^{n}(N_{g})$ admits a spin structure.
This paper is organized as follows. In Section 2, we mention some facts about symmetric products, sequential topological complexity, and sequential distributional complexity. We study the cohomology of symmetric products of finite CW complexes in Section 3, where we give a proof of Theorem C. In Section 4, we prove Theorem D and Corollary D.1. We give a proof of our main result (Theorem A) in Section 5. In this section, we also prove Corollaries A.1A.2, and D.2 and mention several other consequences of our results. We begin Section 6 by proving Theorem E and Corollary E.1. Then we give proofs of our second main result (Theorem B) and Corollary B.1. We end this paper by discussing the problem of determining the topological complexity of symmetric products of non-orientable surfaces.
Notations and conventions
In this paper, the term space refers to a connected CW complex, map refers to a topologically continuous function or a group homomorphism, surface refers to a closed connected $2$-manifold, and ring refers to a commutative ring with unity. We use the symbol “$=$” to denote homeomorphisms of spaces and isomorphisms of groups, and $\simeq$” to denote homotopy equivalences. The tensor product is taken over ${\mathbb{Z}}$ and is denoted by $\otimes$.
2. Preliminaries
In this section, we give a brief review of the symmetric products of spaces and the classical and distributional complexities of spaces and maps.
2.A. Symmetric products
The $n$-th symmetric product of a space $X$, denoted $SP^{n}(X)$, is defined as the orbit space $X^{n}/S_{n}$ of the natural action of the symmetric group $S_{n}$ on the product space $X^{n}$ which permutes the coordinates. Then the elements of $SP^{n}(X)$ are unordered $n$-tuples $[x_{1},\dots,x_{n}]$, where $x_{i}\in X$. We write $\vartheta_{n}:X^{n}\to SP^{n}(X)$ for the corresponding quotient map.
Let us fix some $x_{0}\in X$ as the basepoint of $X$. Then there are the basepoint inclusions $SP^{n}(X)\hookrightarrow SP^{n+k}(X)$ for each $k\geq 1$. The colimit over the symmetric products with respect to the basepoint inclusions defines the infinite symmetric product $SP^{\infty}(X)$, which is a free abelian topological monoid. It follows from the Dold–Thom theorem [DT] that $\pi_{k}(SP^{\infty}(X))=\widetilde{H}_{k}(X;{\mathbb{Z}})$ for each $k\geq 1$.
When $X$ is a CW complex, $SP^{n}(X)$ can also be given a CW structure. Indeed, one can install a CW structure on the product space $X^{n}$ such that each non-identity permutation $\sigma\in S_{n}$ corresponds to a homeomorphism from one simplex of $X^{n}$ to another. Such a CW structure on $X^{n}$ respects the permutation action of the symmetric group $S_{n}$. Hence, this gives a CW complex structure to the orbit space $SP^{n}(X)$, see [tD, Proposition 1.5 and Example 1.17 (2)].
If $X$ is a $2$-dimensional CW complex with only one vertex (for example, if $X$ is a surface), then the CW structure on $SP^{n}(X)$ is natural and can be described explicitly [KS], see also [BR]. More precisely, for some integers $k\geq 0$ and $r\geq 1$, let $X$ be a $2$-dimensional CW complex obtained by attaching $r$ number of closed $2$-discs to the wedge $\bigvee_{i=0}^{k}\hskip 1.42262ptS^{1}_{i}$. Let $e\in X$ be the basepoint of the wedge that is also the identity element of each circle group $S^{1}_{i}$. Define a relation $\sim$ on $SP^{n}(X)$ such that
$\left[x_{1},x_{2},x_{3},\ldots,x_{n}\right]\sim\left[e,x_{1}\cdot x_{2},x_{3},\ldots,x_{n}\right]$
whenever $x_{1},x_{2}\in S^{1}_{i}$ for some $i$. Here, $\cdot$ denotes the group-theoretic multiplication on $S_{i}^{1}$. Let $\overline{SP}^{n}(X):=SP^{n}(X)/\sim$ and $q_{n}:SP^{n}(X)\to\overline{SP}^{n}(X)$ be the quotient map. It follows from [KS] that $\overline{SP}^{n}(X)$ has a canonical CW structure and $q_{n}$ is a natural homotopy equivalence. Also, the composition $q_{n}\circ\vartheta_{n}:X^{n}\to\overline{SP}^{n}(X)$ is cellular. The basepoint inclusions $\overline{SP}^{n}(X)\hookrightarrow\overline{SP}^{n+1}(X)$ define $\overline{SP}^{\infty}(X)$ in the obvious way. Since $q_{n}\circ\vartheta_{n}$ is cellular for each $n$, the $k$-skeleton of $\overline{SP}^{k}(X)$ coincides with the $k$-skeleton of $\overline{SP}^{n}(X)$ for each $k\leq n$. In particular, this means that
(a) the $\text{the }k\text{-skeleton of }\overline{SP}^{\infty}(X)\text{ coincides withthat of }\overline{SP}^{n}(X)\text{ for each }n\geq k.$
If $X$ is a closed manifold (i.e., a compact manifold without boundary), then it is not difficult to show that $SP^{n}(X)$ is a closed manifold for each $n\geq 2$ if and only if $\dim(X)=2$, see [Pu]. For this reason, we will focus more on surfaces and $2$-dimensional CW complexes of the above kind in Sections 5 and 6 of this paper for explicit results and computations.
2.B. LS-category and topological complexity
Let $X$ be a space and $R$ be a ring. Then the cup-length of $H^{*}(X;R)$ (or the $R$-cup-length of $X$), denoted $c\ell_{R}(X)$, is the maximal length $k$ of a non-zero cup product $\alpha_{1}\smile\cdots\smile\alpha_{k}\neq 0$ of positive degree cohomology classes $\alpha_{i}\in H^{*}(X;R)$.
We recall that the Lusternik–Schnirelmann category (LS-category) of a map $f:X\to Y$, denoted $\operatorname{\mathsf{cat}}(f)$, is the smallest integer $n$ for which there is a covering $\{U_{i}\}$ of $X$ by $n+1$ open sets such that the restriction maps $f_{|U_{i}}:U_{i}\to Y$ are null-homotopic for each $i$[BG]. When $f$ is the identity map on $X$, we obtain the LS-category of the space $X$, denoted $\operatorname{\mathsf{cat}}(X)$[Ja][CLOT].
For a proof of the following, see [BG] and [Ja] (see also [Sch] and [CLOT] for the case when $f$ is the identity map).
2.1 Theorem.
For a map $f:X\to Y$ between finite-dimensional CW complexes and ring $R$, $c\ell_{R}(\textup{Im}(f^{*}))\leq\operatorname{\mathsf{cat}}(f)\leq\min\{\operatorname{\mathsf{cat}}(X),\operatorname{\mathsf{cat}}(Y)\}\leq\min\{\dim(X),\dim(Y)\}$.
For any integer $m\geq 2$, let $\Delta_{m}:X\to X^{m}$ denote the diagonal map. Suppose $\Delta_{m}^{*}:H^{*}(X^{m};R)\to H^{*}(X;R)$ is the map induced by $\Delta_{m}$ in $R$-cohomology.
2.2 Definition ([BGRT]).
The $m$-th $R$-zero-divisor cup-length of $X$, denoted $zc\ell^{m}_{R}(X)$, is defined as the cup-length of the ideal $\text{Ker}(\Delta_{m}^{*})$.
For given $m\geq 2$, the $m$-th sequential topological complexity of a map $f:X\to Y$, denoted $\operatorname{\mathsf{TC}}_{m}(f)$, is the smallest integer $n$ for which there exists a covering $\{U_{i}\}$ of $X^{m}$ by $n+1$ open sets over each of which there is a map $s_{i}:U_{i}\to P(Y)$ such that for each $(x_{1},\ldots,x_{m})\in U_{i}\subset X^{m}$,
$s_{i}\left(x_{1},\ldots,x_{m}\right)\left(\frac{j-1}{m-1}\right)=f\left(x_{j}\right)$
for all $1\leq j\leq m$, see [Ku] (and also [Sco] for the case $m=2$). Here, $P(Y)=Y^{[0,1]}$ denotes the path space of $Y$ with the compact-open topology. When $f$ is the identity map on $X$, we obtain the $m$-th sequential topological complexity of the space $X$, denoted $\operatorname{\mathsf{TC}}_{m}(X)$, which was introduced by Rudyak [Ru] as an extension to Farber’s [Far] classical notion of topological complexity $\operatorname{\mathsf{TC}}(X)$.
The next theorem follows from [Sch] and [Ru] (and also [Far] in the case $m=2$).
2.3 Theorem.
For a space $X$, integer $m\geq 2$, and ring $R$, we have the inequalities $\max\{\operatorname{\mathsf{cat}}(X^{m-1}),zc\ell^{m}_{R}(X)\}\leq\operatorname{\mathsf{TC}}_{m}(X)\leq\operatorname{\mathsf{cat}}(X^{m})\leq m\operatorname{\mathsf{cat}}(X)$.
The proof of the following statement, which generalizes Theorem 2.3, can be found in [Ku] (see also [Sco] for the case $m=2$).
2.4 Theorem.
For a map $f:X\to Y$, integer $m\geq 2$, and ring $R$, we have that $\max\{\operatorname{\mathsf{cat}}(f^{m-1}),zc\ell^{m}_{R}(\textup{Im}((f^{m})^{*}))\}\leq\operatorname{\mathsf{TC}}_{m}(f)\leq\min\{\operatorname{\mathsf{cat}}(f^{m}),\operatorname{\mathsf{TC}}_{m}(X),\operatorname{\mathsf{TC}}_{m}(Y)\}$.
Here, $f^{k}:X^{k}\to Y^{k}$ is defined as $f^{k}(x_{1},\ldots,x_{k})=(f(x_{1}),\ldots,f(x_{k}))$ for each $k\geq 1$, and $zc\ell_{R}^{m}(\textup{Im}((f^{m})^{*})):=c\ell_{R}(\textup{Im}((f^{m})^{*})\cap\text{Ker}(\Delta_{m}^{*}))$.
2.C. Distributional category and complexity
Recently in [DJ] and [J1], some probabilistic variants of $\operatorname{\mathsf{cat}}$ and $\operatorname{\mathsf{TC}}_{m}$ were introduced. These are the distributional category and the $m$-th sequential distributional complexity, respectively. Some other closely related probabilistic variants, called the “analog invariants”, were also introduced and studied, independently, by Knudsen and Weinberger in [KW] around the same time. See [DJ, Section 1][KW, Section 1], and [J2, Section 2.A] for some motivations behind these new invariants.
For a metric space $Z$, let ${\mathcal{B}}(Z)$ denote the set of probability measures on $Z$. Also, for any $n\geq 1$, let $\mathcal{B}_{n}(Z)=\{\mu\in\mathcal{B}(Z)\mid|\operatorname{{\rm supp}}(\mu)|\leq n\}$ denote the space of measures on $Z$ supported by at most $n$ points, equipped with the Lévy–Prokhorov metric, [Pr][DJ, Section 3.1]. Let $P(z_{1},z_{2})=\{\phi\in P(Z)\mid\phi(0)=z_{1},\phi(1)=z_{2}\}$ for any $z_{1},z_{2}\in P(Z)$. If $Z$ has a fixed basepoint $z_{0}\in Z$, then we write the based path space of $Z$ as $P_{0}(Z)=\{\phi\in P(Z)\mid\phi(1)=z_{0}\}$.
We recall that the distributional category of a pointed map $f:(X,x_{0})\to(Y,y_{0})$, denoted $\operatorname{\mathsf{dcat}}(f)$, is the smallest integer $n$ for which there exists a continuous map $H:X\to{\mathcal{B}}_{n+1}(P_{0}(Y))$ such that $H(x)(0)\in{\mathcal{B}}_{n+1}(P(f(x),y_{0}))$ for each $x\in X$, see [Dr4][J2]. When $f$ is the identity map on $X$, we obtain the distributional category of the space $X$, denoted $\operatorname{\mathsf{dcat}}(X)$[DJ].
The proof of the following statement is similar to that of Theorem 2.1 (see [DJ] for the case when $f$ is the identity map on $X$).
2.5 Theorem ([Dr4][J2]).
For a map $f:X\to Y$ between finite CW complexes, $c\ell_{{\mathbb{Q}}}(\textup{Im}(f^{*}))\leq\operatorname{\mathsf{dcat}}(f)\leq\min\{\operatorname{\mathsf{cat}}(f),\operatorname{\mathsf{dcat}}(X),\operatorname{\mathsf{dcat}}(Y)\}$.
Let us fix $m\geq 2$. For any $m$-tuple $(z_{1},\ldots,z_{m})\in Z^{m}$, let
$P\left(z_{1},\ldots,z_{m}\right)=\left\{\phi\in P(Z)\hskip 2.84526pt\middle|\hskip 2.84526pt\phi\left(\frac{i-1}{m-1}\right)=z_{i}\textup{ for all }1\leq i\leq m\right\}.$
Given a map $f:X\to Y$ and integer $m\geq 2$, the $m$-th sequential distributional complexity of $f$, denoted $\operatorname{\mathsf{dTC}}_{m}(f)$, is the smallest integer $n$ for which there exists a map $s_{m}:X^{m}\to{\mathcal{B}}_{n+1}(P(Y))$ such that for each $(x_{1},\ldots,x_{m})\in X^{m}$, $s_{m}(x_{1},\ldots,x_{m})\in{\mathcal{B}}_{n+1}(P(f(x_{1}),\ldots,f(x_{m})))$. When $f$ is the identity map on $X$, we obtain the $m$-th sequential distributional complexity of $X$, denoted $\operatorname{\mathsf{dTC}}_{m}(X)$, see [J1] (and also [DJ] for the case $m=2$).
The proof of the following statement is a generalization of the proof of the corresponding statement for $\operatorname{\mathsf{dTC}}_{m}(X)$ found in [J1, Section 3].
2.6 Theorem.
For a map $f:X\to Y$, integer $m\geq 2$, and ring $R$, we have that $\operatorname{\mathsf{dcat}}(f^{m-1})\leq\operatorname{\mathsf{dTC}}_{m}(f)\leq\min\{\operatorname{\mathsf{dcat}}(f^{m}),\operatorname{\mathsf{TC}}_{m}(f),\operatorname{\mathsf{dTC}}_{m}(X),\operatorname{\mathsf{dTC}}_{m}(Y)\}$.
In general, the distributional invariants are different from their classical counterparts. For example, $\operatorname{\mathsf{dcat}}({\mathbb{R}}P^{n})=\operatorname{\mathsf{dTC}}({\mathbb{R}}P^{n})=1<n=\operatorname{\mathsf{cat}}({\mathbb{R}}P^{n})\leq\operatorname{\mathsf{TC}}({\mathbb{R}}P^{n})$ for each $n\geq 2$, see [DJ, Section 3] (and compare with [CLOT, Example 1.8 (2)] and [FTY, Theorem 12]). For various other interesting differences, see [KW][Dr4][J1].
3. Symmetric products of finite CW complexes
In this section, motivated by the problem of lower bounding the LS-category and the sequential topological complexities of the symmetric products of a finite CW complex $X$ in terms of the cohomology of $X$, we prove Theorem C.
Let $X$ be a finite CW complex and $x_{0}\in X$ be a fixed basepoint. Let us write $\xi_{n}:X\to SP^{n}(X)$ for the basepoint inclusion defined as $\xi_{n}(x)=[x,x_{0},\ldots,x_{0}]$, where $x_{0}$ repeats $(n-1)$-times. Then we have the following.
3.1 Theorem ([Na][Do1][Do2]).
For a finite CW complex $X$ and $n\in\mathbb{N}\cup\{\infty\}$, the map $\xi_{n}$ induces a split epimorphism $\xi_{n}^{*}:H^{*}(SP^{n}(X);R)\to H^{*}(X;R)$ in $R$-cohomology for any ring $R$.
Let $\delta_{n}:X\to SP^{n}(X)$ be the diagonal embedding defined as $\delta_{n}(x)=[x,x,\ldots,x]$. Using Theorem 3.1 and the Dold–Thom theorem [DT], the following result was proved in [DJ, Section 4].
3.2 Proposition.
For a finite CW complex $X$ and $n\in\mathbb{N}$, the map $\delta_{n}$ induces a split epimorphism $\delta_{n}^{*}:H^{*}(SP^{n}(X);{\mathbb{Q}})\to H^{*}(X;{\mathbb{Q}})$ in rational cohomology.
See that the proof of Proposition 3.2 from [DJ, Proposition 4.3] also works for cohomology with coefficients in ${\mathbb{R}}$. So, the results of this section that we will prove using Proposition 3.2 in ${\mathbb{Q}}$-coefficients will also be true for ${\mathbb{R}}$-coefficients.
Moving forward, we let $H^{*}(X):=H^{*}(X;{\mathbb{Q}})$.
3.3 Proposition.
For $m\geq 2$, let $X_{j}$ be finite CW complexes for $1\leq j\leq m$. For any $n\geq 2$, let $\delta_{j}:X_{j}\to SP^{n}(X_{j})$ be the diagonal embedding for each $j$. Then the product map $\delta_{1}\times\cdots\times\delta_{m}$ induces a split epimorphism in rational cohomology.
Proof.
Let $\delta_{j}^{i}:H^{i}(SP^{n}(X_{j}))\to H^{i}(X_{j})$ be the map induced by $\delta_{j}$ in the $i$-th rational cohomology. For simplicity, we prove the above statement in the case $m=2$ because the proof for $m\geq 3$ is entirely analogous. We have the following commutative diagram.
(b) ${\bigoplus_{i=0}^{k}\left(H^{i}(X_{1})\otimes H^{k-i}(X_{2})\right)}$${H^{k}\left(X_{1}\times X_{2}\right)}$${\bigoplus_{i=0}^{k}\left(H^{i}(SP^{n}(X_{1}))\otimes H^{k-i}(SP^{n}(X_{2}))\right)}$${H^{k}\left(SP^{n}(X_{1})\times SP^{n}(X_{2})\right).}$$\scriptstyle{\Theta_{1}}$$\scriptstyle{\rho}$$\scriptstyle{\Theta_{2}}$$\scriptstyle{(\delta_{1}\times\delta_{2})^{*}}$
Here, $\Theta_{1}$ and $\Theta_{2}$ are monomorphisms from the short exact sequences in the Kunneth formulas for $X_{1}\times X_{2}$ and $SP^{n}(X_{1})\times SP^{n}(X_{2})$, respectively, [Ha]. Also,
$\rho=\bigoplus_{i=0}^{k}\left(\delta_{1}^{i}\otimes\delta_{2}^{k-i}\right)\hskip 14.22636pt\text{and}\hskip 14.22636pt\Theta_{1}\circ\rho=\left(\delta_{1}\times\delta_{2}\right)^{*}\circ\Theta_{2}.$
Since $X_{j}$ is a finite CW complex for each $j$, $\delta_{j}^{i}$ and $\delta_{j}^{k-i}$ are epimorphisms for each $i$ by Proposition 3.2. Hence, $\rho$ is also an epimorphism. Each $X_{j}$ satisfies the hypotheses of the Kunneth formula for cohomology, so $\Theta_{1}$ is an isomorphism, [Ha, Theorem 3.15]. Therefore, by commutativity of (b), the induced map $(\delta_{1}\times\delta_{2})^{*}$ is an epimorphism. Since all the groups in (b) are finitely generated vector spaces, the map $(\delta_{1}\times\delta_{2})^{*}$ is, in fact, a split epimorphism. This completes the proof. ∎
It is now easy to get lower bounds to $\operatorname{\mathsf{cat}}(SP^{n}(X))$ and $\operatorname{\mathsf{TC}}_{m}(SP^{n}(X))$ in terms of the cohomology of $X$ and $X^{m}$, respectively.
Proof of Theorem C.
Let $c\ell_{R}(X)=r$. Then there exist $\alpha_{i}\in H^{k_{i}}(X;R)$ for some $k_{i}>0$ such that $\alpha_{1}\smile\cdots\smile\alpha_{r}\neq 0$. By Theorem 3.1, there exists a section $\widetilde{s}:H^{*}(X;R)\to H^{*}(SP^{n}(X);R)$ to the map $\xi_{n}^{*}$. Because $\widetilde{s}$ is a monomorphism,
$\widetilde{s}(\alpha_{1})\smile\cdots\smile\widetilde{s}(\alpha_{r})=\widetilde{s}\left(\alpha_{1}\smile\cdots\smile\alpha_{r}\right)\neq 0.$
Hence, $c\ell_{R}(X)=r\leq c\ell_{R}(SP^{n}(X))\leq\operatorname{\mathsf{cat}}(X)$ in view of Theorem 2.1. In the case $R={\mathbb{Q}}$, Proposition 3.2 gives a section $s:H^{*}(X)\to H^{*}(SP^{n}(X))$ to the map $\delta_{n}^{*}$. So, $\alpha_{i}$ has a pre-image under $\delta_{n}^{*}$, namely $s(\alpha_{i})\in H^{k_{i}}(SP^{n}(X))$. Therefore, we have $c\ell_{{\mathbb{Q}}}(X)=r\leq c\ell_{{\mathbb{Q}}}(\text{Im}(\delta_{n}^{*}))$. For any ring $R$, the inequality $c\ell_{R}(\text{Im}(\delta_{n}^{*}))\leq c\ell_{R}(X)$ is straighforward because $\text{Im}(\delta_{n}^{*})$ is a subring of $H^{*}(X;R)$.
Next, let $\Delta_{m}:X\to X^{m}$ and $\partial_{m}:SP^{n}(X)\to(SP^{n}(X))^{m}$ denote the diagonal inclusions. Let $zc\ell^{m}_{{\mathbb{Q}}}(X)=r$. Then there exist $\beta_{i}\in H^{k_{i}}(X^{m})$ for some $k_{i}>0$ such that $\Delta_{m}^{*}(\beta_{i})=0$ for each $1\leq i\leq r$ and $\beta_{1}\smile\cdots\smile\beta_{r}\neq 0$. Consider the following commutative diagram
(c) ${H^{*}\left((SP^{n}(X))^{m}\right)}$${H^{*}(X^{m})}$${H^{*}\left(SP^{n}(X)\right)}$${H^{*}(X),}$$\scriptstyle{(\delta_{n}^{m})^{*}}$$\scriptstyle{\partial_{m}^{*}}$$\scriptstyle{{s_{m}\vphantom{1}}}$$\scriptstyle{\Delta_{m}^{*}}$$\scriptstyle{\delta_{n}^{*}}$$\scriptstyle{{\overline{s}\vphantom{1}}}$
where $\overline{s}:H^{*}(X)\to H^{*}(SP^{n}(X))$ and $s_{m}:H^{*}(X^{m})\to H^{*}((SP^{n}(X))^{m})$ are sections to the maps $\delta_{n}^{*}$ and $(\delta_{n}^{m})^{*}$, respectively, implied by Propositions 3.2 and 3.3. By commutativity in (c), we have for each $1\leq i\leq r$ that
$\partial_{m}^{*}(s_{m}(\beta_{i}))=\overline{s}(\Delta_{m}^{*}(\beta_{i}))=0.$
Hence, $s_{m}(\beta_{i})$ is a $m$-th rational zero-divisor of $SP^{n}(X)$ for each $i$. As a result, we get $zc\ell^{m}_{{\mathbb{Q}}}(X)=r\leq zc\ell^{m}_{{\mathbb{Q}}}(SP^{n}(X))\leq\operatorname{\mathsf{TC}}_{m}(SP^{n}(X))$ by Theorem 2.3. Note that $s_{m}(\beta_{i})\in H^{k_{i}}((SP^{n}(X))^{m})$ is the preimage of $\beta_{i}$ under $(\delta_{n}^{m})^{*}$. Therefore, we have $zc\ell^{m}_{{\mathbb{Q}}}(X)=r\leq zc\ell^{m}_{{\mathbb{Q}}}(\text{Im}((\delta_{n}^{m})^{*}))$. The reverse inequality is obvious. ∎
Theorems C2.1, and 2.4 imply the following.
3.4 Corollary.
For any finite CW complex $X$ and integers $m\geq 2$ and $n\geq 1$, $c\ell_{{\mathbb{Q}}}(X)\leq\operatorname{\mathsf{cat}}(\delta_{n})$ and $zc\ell_{{\mathbb{Q}}}^{m}(X)\leq\operatorname{\mathsf{TC}}_{m}(\delta_{n})$.
We conclude this section by making some simple observations.
3.5 Remark.
For each $n\geq 1$, it is known that $SP^{n}(S^{1})\simeq S^{1}$ (see [KS]). Hence, $\operatorname{\mathsf{TC}}_{m}(SP^{n}(S^{1}))=zc\ell_{{\mathbb{Q}}}^{m}(S^{1})=m-1$ (see [Ru]) and $\operatorname{\mathsf{cat}}(SP^{n}(S^{1}))=c\ell_{{\mathbb{Q}}}(S^{1})=1$. Thus, the lower bounds obtained in Theorem C and Corollary 3.4 are sharp.
3.6 Example.
Due to Corollary 3.4, $\operatorname{\mathsf{TC}}_{m}(\delta_{n})$ (resp. $\operatorname{\mathsf{cat}}(\delta_{n})$) is determined, to be equal to $\operatorname{\mathsf{TC}}_{m}(X)$ (resp. $\operatorname{\mathsf{cat}}(X)$), whenever $zc\ell^{m}_{{\mathbb{Q}}}(X)=\operatorname{\mathsf{TC}}_{m}(X)$ (resp. $c\ell_{{\mathbb{Q}}}(X)=\operatorname{\mathsf{cat}}(X)$). In particular, this holds when $X$ is a simply connected symplectic manifold, a finite product of spheres, or a symmetric product of orientable surfaces. For the first two classes, see [BGRT, Section 3]; for the last class, see proofs of Theorems A and 5.3.
4. Sequential distributional complexity
A lower bound to the $m$-th sequential distributional complexity, $\operatorname{\mathsf{dTC}}_{m}(X)$, for each $m\geq 2$ was obtained in [J1, Theorem 4.4] in terms of the Alexander–Spanier cohomology of $SP^{k}(X^{m})$. In this section, we will prove Theorem D, which gives a lower bound to $\operatorname{\mathsf{dTC}}_{m}(X)$ in terms of the $m$-th zero-divisors of $SP^{k}(X)$.
Let us fix some integers $m,n\geq 2$. For each $1\leq i\leq m$, given the projection maps $\pi_{i}:Y^{m}\to Y$, let $r_{i}=SP^{n!}(\pi_{i}):SP^{n!}(Y^{m})\to SP^{n!}(Y)$ be the induced maps. Define $r:SP^{n!}(Y^{m})\to(SP^{n!}(Y))^{m}$ as the map $r=(r_{1},\ldots,r_{m})$. Suppose that $\delta_{n!}:Y\to SP^{n!}(Y)$ and $\partial_{m}:Y^{m}\to SP^{n!}(Y^{m})$ are the respective diagonal inclusions. Then, $r\circ\partial_{m}=\delta_{n!}^{m}$ for the product map $\delta_{n!}^{m}:Y^{m}\to(SP^{n!}(Y))^{m}$.
Let us fix a map $f:X\to Y$ and consider the following two pullback diagrams.
[…]
Here, $\zeta_{n}=SP^{n!}(\tau_{m})$, where $\tau_{m}:P(Y)\to Y^{m}$ is the fibration defined as
$\tau_{m}(\phi)=\left(\phi(0),\phi\left(\frac{1}{m-1}\right),\ldots,\phi\left(\frac{m-2}{m-1}\right),\phi(1)\right).$
We will use the same notations in the above pullback diagrams when the space $X^{m}$ is replaced with any subspace $A_{i}\subset X^{m}$.
4.1 Lemma.
Given a map $f:X\to Y$ and integer $m\geq 2$, if $\operatorname{\mathsf{dTC}}_{m}(f)<n$, then there exist sets $A_{1},A_{2},\ldots,A_{n}$ that cover $X^{m}$ such that over each $A_{i}$ there is a section to the map $\theta_{n}$ from (d).
Proof.
Suppose that $\operatorname{\mathsf{dTC}}_{m}(f)<n$. Then there exists a map $s_{m}:X^{m}\to{\mathcal{B}}_{n}(P(Y))$ as defined in Section 2.C. Proceeding as in [J1, Lemma 4.3], we define sets
$A_{i}=\left\{\overline{x}=(x_{1},\ldots,x_{m})\in X^{m}\mid|\operatorname{{\rm supp}}(s_{m}(\overline{x}))|=i\right\}$
for each $1\leq i\leq n$ and corresponding maps $H_{i}:A_{i}\to SP^{n!}(P(Y))$ such that
$H_{i}(\overline{x})=\sum_{\phi\hskip 2.84526pt\in\hskip 2.84526pt\operatorname{{\rm supp}}(s_{m}(\overline{x}))}\hskip 2.84526pt\frac{n!}{i}\phi.$
Clearly, $\{A_{i}\}_{i=1}^{n}$ is a cover of $X^{m}$ and $\zeta_{n}\circ H_{i}=\partial_{m}f^{m}$ for each $1\leq i\leq n$, where we use $f^{m}$ to denote the restriction of $f^{m}$ to $A_{i}$. Then, $(r\circ\zeta_{n})\circ H_{i}=\delta_{n!}^{m}\circ f^{m}$. So, by the universal property of the pullback on the right in (d), there exists a map $K_{i}:A_{i}\to\mathcal{Z}_{n,m}$ such that $\theta_{n}\circ K_{i}$ is the identity map on $A_{i}$. ∎
Proof of Theorem D.
Let $\Delta^{n!}:SP^{n!}(Y)\to(SP^{n!}(Y))^{m}$ be the diagonal map. Define a homotopy equivalence $g:SP^{n!}(Y)\to SP^{n!}(P(Y))$ as
$g\left[a_{1},\ldots,a_{n!}\right]=\left[c_{a_{1}},\ldots,c_{a_{n!}}\right],$
where $c_{a_{j}}$ denotes the trivial loop at $a_{j}\in Y$, see [J1, Lemma 4.2]. If we assume that $\operatorname{\mathsf{dTC}}_{m}(f)<n$, then by Lemma 4.1, we get sets $A_{i}\subset X^{m}$ and maps $K_{i}:A_{i}\to\mathcal{Z}_{n,m}$ such that $\bigcup_{i=1}^{n}A_{i}=X^{m}$ and $\theta_{n}\circ K_{i}=\mathbbm{1}_{A_{i}}$. For each fixed $1\leq i\leq n$, we have the following commutative diagram in $R$-cohomology.
[…]
By hypothesis, $(\Delta^{n!})^{*}(\alpha_{i}^{*})=0$. Because $g^{*}$ is an isomorphism, $(r\circ\zeta_{n})^{*}(\alpha_{i}^{*})=0$. From the commutative square in (e), we get that
$(\delta_{n!}^{m}\circ f^{m})^{*}(\alpha_{i}^{*})=(K_{i}^{*}\circ b^{*})((r\circ\zeta_{n})^{*}(\alpha_{i}^{*}))=0.$
Seeing the map $\delta_{n!}^{m}\circ f^{m}:A_{i}\to(SP^{n!}(Y))^{m}$ as the inclusion, we have the following long exact sequence of the pair $((SP^{n!}(Y))^{m},A_{i})$ in Alexander–Spanier cohomology:
$\cdots\to H^{k_{i}}((SP^{n!}(Y))^{m},A_{i};R)\xrightarrow{j_{i}^{*}}H^{k_{i}}((SP^{n!}(Y))^{m};R)\xrightarrow{(\delta_{n}^{m}f^{m})^{*}}H^{k_{i}}(A_{i};R)\to\cdots,$
where $j_{i}:(SP^{n!}(Y))^{m}\hookrightarrow((SP^{n!}(Y))^{m},A_{i})$. Since $(\delta_{n!}^{m}\circ f^{m})^{*}(\alpha_{i}^{*})=0$, there exists a class $\overline{\alpha_{i}^{*}}\in H^{k_{i}}((SP^{n!}(Y))^{m}),A_{i};R)$ such that $j_{i}^{*}(\overline{\alpha_{i}^{*}})=\alpha_{i}^{*}$. Also, let $(\delta_{n!}^{m}\circ f^{m})^{*}(\overline{\alpha_{i}^{*}})=\overline{\alpha_{i}}\in H^{k_{i}}(X^{m},A_{i};R)$. For the maps $j=\sum j_{i}$ and $j^{\prime}=\sum j_{i}^{\prime}$, where $j_{i}^{\prime}:X^{m}\hookrightarrow(X^{m},A_{i})$, and integer $k=\sum k_{i}$, we get the following commutative diagram in Alexander–Spanier cohomology.
(f) ${H^{k}(X^{m};R)}$${H^{k}\left(X^{m},\bigcup_{i=1}^{n}A_{i};R\right)}$${H^{k}\left((SP^{n!}(Y))^{m};R\right)}$${H^{k}\left((SP^{n!}(Y))^{m},\bigcup_{i=1}^{n}A_{i};R\right).}$$\scriptstyle{(j^{\prime})^{*}}$$\scriptstyle{(\delta^{m}_{n!}f^{m})^{*}}$$\scriptstyle{(\delta^{m}_{n!}f^{m})^{*}}$$\scriptstyle{j^{*}}$
The cup product $\overline{\alpha_{1}^{*}}\smile\cdots\smile\overline{\alpha_{n}^{*}}$ in the bottom-right part of (f) goes to the non-zero cup product $\alpha_{1}\smile\cdots\smile\alpha_{n}\neq 0$ in the top-left part part of (f). But in the process, it factors through $\overline{\alpha_{1}}\smile\cdots\smile\overline{\alpha_{n}}\in H^{k}(X^{m},X^{m};R)=0$. This is a contradiction. Therefore, we must have $\operatorname{\mathsf{dTC}}_{m}(X)\geq n$. ∎
Basically, for the diagonal map $\Delta^{n!}:SP^{n!}(Y)\to(SP^{n!}(Y))^{m}$ and the map $\delta_{n!}^{m}\circ f^{m}:X^{m}\to(SP^{n!}(Y))^{m}$, we have for each fixed $m\geq 2$ and ring $R$ that
$c\ell_{R}(\textup{Im}((\delta_{n!}^{m}\circ f^{m})^{*}:\textup{Ker}((\Delta^{n!})^{*})\to H^{*}(X^{m};R)))\geq n\implies\operatorname{\mathsf{dTC}}_{m}(f)\geq n.$
Proof of Corollary D.1.
Let $zc\ell^{m}_{{\mathbb{Q}}}(X)=n$. Then there are $m$-th rational zero-divisors of $X$, say $u_{i}\in H^{*}(X^{m};{\mathbb{Q}})$, such that $u_{1}\smile\cdots\smile u_{n}\neq 0$. By the proof of Theorem C, for each $i$, there exists an $m$-th rational zero-divisor of $SP^{n!}(X)$, say $v_{i}\in H^{*}((SP^{n!}(X))^{m};{\mathbb{Q}})$, such that $(\delta_{n!}^{m})^{*}(v_{i})=u_{i}$. On finite CW complexes, singular cohomology and Alexander–Spanier cohomology coincide, [Sp]. Hence, we can apply Theorem D, by taking $f$ to be the identity map on $X$ and $R={\mathbb{Q}}$, with singular cohomology classes $v_{i}$ for $1\leq i\leq n$ to get $\operatorname{\mathsf{dTC}}_{m}(X)\geq n=zc\ell_{{\mathbb{Q}}}^{m}(X)$. ∎
5. Symmetric products of orientable surfaces
Let $M_{g}$ denote the orientable surface of genus $g\geq 0$. In this section, we shall prove Theorem A, which computes the $m$-th sequential topological complexity of the symmetric $n$-th product $SP^{n}(M_{g})$ for each $m,n\geq 2$ and $g\geq 0$ using the $m$-th rational zero-divisors of $SP^{n}(M_{g})$.
From now onwards, for brevity, we skip the symbol “ $\smile$ ” while writing cup products, i.e., we let $xy:=x\smile y$.
5.A. Cohomology
Throughout this section, we let $H^{*}(X):=H^{*}(X;{\mathbb{Z}})$ and $H_{*}(X):=H_{*}(X;{\mathbb{Z}})$.
The orientable surface $M_{g}$ can be viewed as a connected sum of $g$ copies of $2$-tori, say ${\mathbb{T}}_{i}$ for $1\leq i\leq g$. The CW structure on $M_{g}$ consists of one 00-cell, $2g$ $1$-cells, and one $2$-cell attached by the product of commutators $[a_{1},b_{1}]\cdots[a_{g},b_{g}]$. Here, $a_{i},b_{i}\in H_{1}({\mathbb{T}}_{i})$ are the generators for each $i$. Let us write $J:=\prod_{i=1}^{g}{\mathbb{T}}_{i}$ and $\text{proj}_{i}:J\to{\mathbb{T}}_{i}$ for the projection map. Let $\mu_{n}:SP^{n}(M_{g})\to J$ be the Abel–Jacobi map (see [DDJ, Sections 2.C & 2.D] for its description).
Let $a_{i}^{*}$ and $b_{i}^{*}$ denote the Hom duals of $a_{i}$ and $b_{i}$, respectively. We use the same notations for their images in $H^{1}(SP^{n}(M_{g}))$ under the induced map $(\text{proj}_{i}\circ\mu_{n})^{*}$. Let $c\in H_{2}(M_{g})$ denote the fundamental class and its image in $H_{2}(SP^{n}(M_{g}))$ under the map induced in homology by the basepoint inclusion $M_{g}\hookrightarrow SP^{n}(M_{g})$. We denote the Hom dual of $c\in H_{2}(SP^{n}(M_{g}))$ by $c^{*}$. Note that in $H^{*}(M_{g})$, we have $a_{i}^{*}b_{j}^{*}=a_{i}^{*}a_{j}^{*}=b_{i}^{*}b_{j}^{*}=0$ for each $i\neq j$ and $(a_{i}^{*})^{2}=(b_{i}^{*})^{2}=0$[Ha]. Also, $c^{*}$ commutes with $a_{i}^{*}$ and $b_{j}^{*}$, and $a_{i}^{*}$ and $b_{j}^{*}$ anti-commute with each other in $H^{*}(SP^{n}(M_{g}))$.
The following famous theorem is due to Macdonald, [Mac].
5.1 Theorem.
The integral cohomology ring $H^{*}(SP^{n}(M_{g}))$ is the quotient of $\Lambda_{{\mathbb{Z}}}(a_{1}^{*},b_{1}^{*},\dots,a_{g}^{*},b_{g}^{*})\otimes{\mathbb{Z}}[c^{*}]$ by the following relation:
$a_{i_{1}}^{*}\cdots a_{i_{\ell}}^{*}b_{j_{1}}^{*}\cdots b_{j_{m}}^{*}(c^{*}-a_{k_{1}}^{*}b_{k_{1}}^{*})\cdots(c^{*}-a_{k_{r}}^{*}b_{k_{r}}^{*})(c^{*})^{s}=0$
whenever $\ell+m+2r+s\geq n+1$ for any collection of distinct indices $i_{1},\dots,i_{\ell}$, $j_{1},\dots,j_{m}$, and $k_{1},\dots,k_{r}$.
The following statement was proven in [DDJ, Proposition 3.3] using the above description.
5.2 Proposition.
In the integral cohomology ring $H^{*}(SP^{n}(M_{g}))$,
  1. (1)
    if $n\leq g$, then the cup product $a_{1}^{*}b_{1}^{*}\cdots a_{n}^{*}b_{n}^{*}$ is non-zero, and
  2. (2)
    if $n>g$, then the cup product $a_{1}^{*}b_{1}^{*}\cdots a_{g}^{*}b_{g}^{*}(c^{*})^{n-g}$ is non-zero.
The integral cohomology ring $H^{*}(SP^{n}(M_{g}))$ is torsion-free, [Mac][KS]. So, $a_{i}^{*}$, $b_{j}^{*}$, $c^{*}\in H^{*}(SP^{n}(M_{g}))$ can be considered as rational cohomology classes and hence, Theorem 5.1 and Proposition 5.2 hold true in rational cohomology as well.
The LS-category of $SP^{n}(M_{g})$ can now be computed easily.
5.3 Theorem ([DDJ, Theorem 6.5]).
For each $n\geq 1$ and $g\geq 0$, we have that
$\operatorname{\mathsf{cat}}\left(SP^{n}(M_{g})\right)=\begin{cases}2n&\text{if}n\leq g\\ n+g&\text{if }n>g.\end{cases}$
Proof.
For $n\leq g$, we have due to Proposition 5.2 (1) and Theorem 2.1 that
$2n\leq c\ell_{{\mathbb{Q}}}(SP^{n}(M_{g}))\leq\operatorname{\mathsf{cat}}(SP^{n}(M_{g}))\leq\dim(SP^{n}(M_{g}))=2n.$
Similarly, for $n>g$, Proposition 5.2 (2) and Theorem 2.1 imply the inequality $n+g\leq c\ell_{{\mathbb{Q}}}(SP^{n}(M_{g}))\leq\operatorname{\mathsf{cat}}(SP^{n}(M_{g}))$. In this case, the main result of [Dr3] gives
$\operatorname{\mathsf{cat}}(SP^{n}(M_{g}))\leq\frac{\dim(SP^{n}(M_{g}))+\operatorname{{\rm cd}}(\pi_{1}(SP^{n}(M_{g})))}{2}=\frac{2n+2g}{2}=n+g.$
This completes the proof. ∎
5.B. Sequential topological complexity
For $m\geq 2$, let $\Delta_{m}:X\to X^{m}$ denote the diagonal map. For each $1\leq i\leq m$, let $\pi_{i}:X^{m}\to X$ be the projection onto the $i$-th coordinate. For any $y\in H^{*}(X;{\mathbb{Q}})$, let us define $\overline{y}\in H^{*}(X^{m};{\mathbb{Q}})$ as
$\overline{y}=\sum_{i=1}^{m-1}\pi_{i}^{*}(y)-(m-1)\pi_{m}^{*}(y).$
Here, $\pi_{i}^{*}(y)=1\otimes\cdots\otimes 1\otimes y\otimes 1\otimes\cdots\otimes 1\in H^{*}(X^{m};{\mathbb{Q}})$ for each $1\leq i\leq m$, where $y$ appears in the $i$-th position. We note that $\Delta_{m}^{*}(\overline{y})=(m-1)y-(m-1)y=0$, see [Ru, Section 4]. Hence, $\overline{y}$ is a $m$-th rational zero-divisor of $X$.
We will follow these notations in the proof of our main result.
Proof of Theorem A.
Let us fix $m\geq 2$. For each $n\geq 2$ and $g\geq 0$, we have the upper bound $\operatorname{\mathsf{TC}}_{m}(SP^{n}(M_{g}))\leq m\operatorname{\mathsf{cat}}(SP^{n}(M_{g}))$ due to Theorem 2.3, where the value $\operatorname{\mathsf{cat}}(SP^{n}(M_{g}))$ is known by Theorem 5.3. Let us now find the lower bounds separately in the cases $n\leq g$ and $n>g$.
We note that the classes $\overline{a_{i}^{*}},\overline{b_{j}^{*}},\overline{c^{*}}\in H^{*}((SP^{n}(M_{g}))^{m};{\mathbb{Q}})$ are non-zero. We also recall that $\overline{a_{i}^{*}},\overline{b_{j}^{*}},\overline{c^{*}}\in\text{Ker}(\Delta_{m}^{*})$ for each $i,j$. For each $k\geq 2$, $(a_{i}^{*})^{k}=(b_{j}^{*})^{k}=0$. Hence, it follows from the relationship between the cup product and the tensor product [Ha, Section 3.2] that for each $1\leq i,j\leq g$,
[…]
Let $n\leq g$. Then $a_{1}^{*}b_{1}^{*}\cdots a_{n}^{*}b_{n}^{*}\neq 0$ by Proposition 5.2 (1). Hence, the following cup product is non-zero:
$(\overline{a_{1}^{*}})^{m}(\overline{b_{1}^{*}})^{m}\cdots(\overline{a_{n}^{*}})^{m}(\overline{b_{n}^{*}})^{m}=\left((m-1)m!\right)^{2n}\underbrace{\left(\left(a_{1}^{*}b_{1}^{*}\cdots a_{n}^{*}b_{n}^{*}\right)\otimes\cdots\otimes\left(a_{1}^{*}b_{1}^{*}\cdots a_{n}^{*}b_{n}^{*}\right)\right)}_{\text{$m$-times}}.$
Therefore, we get $2mn\leq zc\ell_{{\mathbb{Q}}}^{m}(SP^{n}(M_{g}))\leq\operatorname{\mathsf{TC}}_{m}(SP^{n}(M_{g}))$ from Theorem 2.3. Thus, we have that $\operatorname{\mathsf{TC}}_{m}(SP^{n}(M_{g}))=2mn$ in the case $n\leq g$.
Next, we let $n>g$. It is not difficult to see that $(\overline{c^{*}})^{m(n-g)}$ contains the term
$(m-1)\prod_{i=0}^{m-1}{(m-i)(n-g)\choose n-g}\underbrace{\left((c^{*})^{n-g}\otimes\cdots\otimes(c^{*})^{n-g}\right)}_{\text{$m$-times}}.$
By Proposition 5.2 (2), $a_{1}^{*}b_{1}^{*}\cdots a_{g}^{*}b_{g}^{*}(c^{*})^{n-g}\neq 0$. Hence, the above term and the following term are both non-zero.
$(\overline{a_{1}^{*}})^{m}(\overline{b_{1}^{*}})^{m}\cdots(\overline{a_{g}^{*}})^{m}(\overline{b_{g}^{*}})^{m}=\left((m-1)m!\right)^{2g}\underbrace{\left(\left(a_{1}^{*}b_{1}^{*}\cdots a_{g}^{*}b_{g}^{*}\right)\otimes\cdots\otimes\left(a_{1}^{*}b_{1}^{*}\cdots a_{g}^{*}b_{g}^{*}\right)\right)}_{\text{$m$-times}}.$
Also, for each $k\geq 1$, it follows from Theorem 5.1 that $a_{1}^{*}b_{1}^{*}\cdots a_{g}^{*}b_{g}^{*}(c^{*})^{n-g+k}=0$. Thus, we finally have that the cup product $(\overline{a_{1}^{*}})^{m}(\overline{b_{1}^{*}})^{m}\cdots(\overline{a_{g}^{*}})^{m}(\overline{b_{g}^{*}})^{m}(\overline{c^{*}})^{m(n-g)}$ is equal to the following non-zero term:
$(m-1)^{2g+1}(m!)^{2g}\prod_{i=0}^{m-1}{(m-i)(n-g)\choose n-g}\underbrace{\left(\left(a_{1}^{*}b_{1}^{*}\cdots a_{g}^{*}b_{g}^{*}(c^{*})^{n-g}\right)\otimes\cdots\otimes\left(a_{1}^{*}b_{1}^{*}\cdots a_{g}^{*}b_{g}^{*}(c^{*})^{n-g}\right)\right)}_{\text{$m$-times}}.$
Therefore, $2mg+m(n-g)=m(n+g)\leq zc\ell_{{\mathbb{Q}}}^{m}(SP^{n}(M_{g}))\leq\operatorname{\mathsf{TC}}_{m}(SP^{n}(M_{g}))$ due to Theorem 2.3. Hence, $\operatorname{\mathsf{TC}}_{m}(SP^{n}(M_{g}))=mn+mg$ in the case $n>g$. ∎
For completeness, we recall that for any integer $m\geq 2$, $\operatorname{\mathsf{TC}}_{m}(M_{0})=m$, $\operatorname{\mathsf{TC}}_{m}(M_{1})=2m-2$, and $\operatorname{\mathsf{TC}}_{m}(M_{g})=2m$ for each $g\geq 2$.
We now verify the rationality conjecture (Conjecture 1.1) for $SP^{n}(M_{g})$ for each $n\geq 2$ and $g\geq 0$. We note that this conjecture has been verified for only a few classes of spaces, namely the spheres, simply connected symplectic manifolds, compact Lie groups, surfaces (ii)(ii)(ii) The verification of the rationality conjecture for non-orientable surfaces seems to be missing from [FO] and the literature thereafter. Indeed, Conjecture 1.1 holds for $N_{g}$ in view of the computations of [GGGL][Dr2], and [CV]. The rational function represented by the power series of the $\operatorname{\mathsf{TC}}$-generating function of $N_{g}$ in the case $g=1$ is different from that in the cases $g\geq 2$., and the classifying spaces of Higman’s group, Right-angled Artin groups, and torsion-free hyperbolic groups, see [FO] and [HL].
Proof of Corollary A.1.
Using Theorem A, we see that the $\operatorname{\mathsf{TC}}$-generating function of $X=SP^{n}(M_{g})$ is the power series
$f_{SP^{n}(M_{g})}(t)=\sum_{m=1}^{\infty}(m+1)\operatorname{\mathsf{cat}}(SP^{n}(M_{g}))\hskip 1.42262ptt^{m}=\operatorname{\mathsf{cat}}(SP^{n}(M_{g}))\hskip 1.42262pt\frac{d}{dt}\left(\sum_{m=1}^{\infty}t^{m+1}\right),$
which represents the rational function
$h_{SP^{n}(M_{g})}(t)=\frac{(2t-t^{2})\operatorname{\mathsf{cat}}(SP^{n}(M_{g}))}{(1-t)^{2}}.$
Clearly, the numerator of $h_{SP^{n}(M_{g})}$ is an integral quadratic polynomial whose value at $t=1$ is $\operatorname{\mathsf{cat}}(SP^{n}(M_{g}))$. ∎
Proof of Corollary A.2.
This follows directly from Theorems A and 2.3. ∎
We now show that the sequential distributional complexities of $SP^{n}(M_{g})$ and the distributional category of its products agree with their classical counterparts.
Proof of Corollary D.2.
For each $m\geq 2$, the proof of Theorem A gives the equality $\operatorname{\mathsf{TC}}_{m}(SP^{n}(M_{g}))=zc\ell_{{\mathbb{Q}}}^{m}(SP^{n}(M_{g}))$. Hence, Corollary D.1 and Theorem 2.6 imply that $\operatorname{\mathsf{dTC}}_{m}(SP^{n}(M_{g}))=\operatorname{\mathsf{TC}}_{m}(SP^{n}(M_{g}))$. Using this and Corollary A.2, we get from Theorems 2.5 and 2.6 that $\operatorname{\mathsf{dcat}}((SP^{n}(M_{g}))^{m})=\operatorname{\mathsf{cat}}((SP^{n}(M_{g}))^{m})$. ∎
5.C. More consequences
Let us highlight some ways in which the LS-category and the sequential topological complexity of $SP^{n}(M_{g})$ are “nicely behaved”.
  1. (1)
    Logarithmic law conjecture: For spaces $Y_{i}$[Ja, Proposition 2.3] implies that
    $\operatorname{\mathsf{cat}}\left(Y_{1}\times Y_{2}\times\cdots\times Y_{m}\right)\leq\sum_{i=1}^{m}\operatorname{\mathsf{cat}}\left(Y_{i}\right)$
    for any $m\geq 1$. This leads to the “logarithmic law conjecture”, which says that for any $m\geq 1$ and closed manifold $Y$, $\operatorname{\mathsf{cat}}(Y^{m})=m\operatorname{\mathsf{cat}}(Y)$. This conjecture is not true in general, see [Dr1] and the references therein. However, for $Y=SP^{n}(M_{g})$, this conjecture is verified for each $m,n$, and $g$ in view of Corollary A.2.
  2. (2)
    The Ganea conjecture: For any space $Y$, $\operatorname{\mathsf{cat}}(Y\times S^{1})\leq\operatorname{\mathsf{cat}}(Y)+1$. Ganea’s conjecture [CLOT] says that the above inequality should be an equality, in particular, if $Y$ is a closed manifold. This conjecture was disproved by Iwase in [Iw]. Interestingly, we can verify this conjecture for $Y=SP^{n}(M_{g})$ for each $n$ and $g$. Using techniques from the proof of Theorem 5.3, it can be shown that
    $\operatorname{\mathsf{cat}}(SP^{n}(M_{g})\times S^{1})=\begin{cases}2n+1&\text{ if }n\leq g\\ n+g+1&\text{ if }n>g.\end{cases}$
  3. (3)
    Growth of the sequence $\{\operatorname{\mathsf{TC}}_{m}(X)\}_{m\geq 2}$: For a space $X$ and $m\geq 2$, we have from [BGRT, Proposition 3.7] that $\operatorname{\mathsf{TC}}_{m+1}(X)-\operatorname{\mathsf{TC}}_{m}(X)\leq 2\operatorname{\mathsf{cat}}(X)$. Here, for each $m\geq k$,
    $\operatorname{\mathsf{TC}}_{m}(SP^{n}(M_{g}))-\operatorname{\mathsf{TC}}_{k}(SP^{n}(M_{g}))=(m-k)\operatorname{\mathsf{cat}}(SP^{n}(M_{g}))$
    due to Theorem A. In particular, the sequence $\{\operatorname{\mathsf{TC}}_{m}(SP^{n}(M_{g}))\}_{m\geq 2}$ is an arithmetic progression with common difference $\operatorname{\mathsf{cat}}(SP^{n}(M_{g}))$.
Let us now make more observations using Theorems A and 5.3.
  1. (i)
    Let $G$ be a compact Lie group acting on a finite CW complex $X$. Let $\operatorname{\mathsf{cat}}_{G}(X)$ and $\operatorname{\mathsf{TC}}_{G}(X)$ denote the $G$-equivariant LS-category [Mar] and the $G$-equivariant topological complexity [CG] of $X$, respectively. While we have
    $\max\{\operatorname{\mathsf{cat}}(X/G),\operatorname{\mathsf{cat}}(X)\}\leq\operatorname{\mathsf{cat}}_{G}(X)\hskip 9.95845pt\text{and}\hskip 8.53581pt\max\{\operatorname{\mathsf{TC}}(X/G),\operatorname{\mathsf{TC}}(X)\}\leq\operatorname{\mathsf{TC}}_{G}(X),$
    there is no relation between $\operatorname{\mathsf{cat}}(X/G)$ and $\operatorname{\mathsf{cat}}(X)$ and between $\operatorname{\mathsf{TC}}_{m}(X/G)$ and $\operatorname{\mathsf{TC}}_{m}(X)$ for any $m\geq 2$. To see this, first, take $G={\mathbb{Z}}_{2}$ and $X=S^{n}$ for $n\geq 2$. Note that $X/G={\mathbb{R}}P^{n}$. Clearly, $\operatorname{\mathsf{cat}}(X)=1<n=\operatorname{\mathsf{cat}}(X/G)$ and $\operatorname{\mathsf{TC}}_{m}(X)\leq m<\operatorname{\mathsf{TC}}_{m}(X/G)$. Next, let us take $G=S_{n}$ and $X=(M_{g})^{n}$ for $n>g\geq 2$. Note that $X/G=SP^{n}(M_{g})$. By Theorems 5.3 and A, we get $\operatorname{\mathsf{cat}}(X/G)=n+g<2n=\operatorname{\mathsf{cat}}(X)$ and $\operatorname{\mathsf{TC}}_{m}(X/G)=mn+mg<\operatorname{\mathsf{TC}}_{m}(X)$.
  2. (ii)
    It follows from [BGRT, Proposition 3.11] that $\operatorname{\mathsf{TC}}_{m}(X^{n})\leq n\operatorname{\mathsf{TC}}_{m}(X)$ for any $m\geq 2$ and $n\geq 1$. However, it turns out that in general, there is no relation between $\operatorname{\mathsf{TC}}_{m}(SP^{n}(X))$ and $n\operatorname{\mathsf{TC}}_{m}(X)$. For each $n>g\geq 2$, the inequality $\operatorname{\mathsf{TC}}_{m}(SP^{n}(M_{g}))<n\operatorname{\mathsf{TC}}_{m}(M_{g})$ follows from the above item. Using Theorem A, we also see that $\operatorname{\mathsf{TC}}(SP^{n}({\mathbb{T}}))>n\operatorname{\mathsf{TC}}({\mathbb{T}})\geq\operatorname{\mathsf{TC}}({\mathbb{T}}^{n})$ for each $n\geq 2$ and that $\operatorname{\mathsf{TC}}_{3}(SP^{2}({\mathbb{T}}))>2\operatorname{\mathsf{TC}}_{3}({\mathbb{T}})\geq\operatorname{\mathsf{TC}}_{3}({\mathbb{T}}^{2})$ for the $2$-torus ${\mathbb{T}}$.
  3. (iii)
    It is well-known that $SP^{2}(M_{1})$ is a $S^{2}$-bundle over a $2$-torus, say ${\mathbb{T}}$. Let $p:SP^{2}(M_{1})\to{\mathbb{T}}$ be the corresponding locally trivial fibration with fiber $S^{2}$. Then for the $m$-th sequential parametrized topological complexity of the fibration $p$ (see [FP], and also [CFW] for the case $m=2$), we have because of [FP, Definition 3.1 and Proposition 6.1] that
    $m=\operatorname{\mathsf{TC}}_{m}\left(S^{2}\right)\leq\operatorname{\mathsf{TC}}_{m}\left[\hskip 1.13809ptp:SP^{2}\left(M_{1}\right)\to{\mathbb{T}}\right]\leq\frac{2m+2+1}{2}=m+\frac{3}{2}.$
    However, $\operatorname{\mathsf{TC}}_{m}(SP^{2}(M_{1}))=3m$ by Theorem A. Hence, for each $m\geq 2$, $\operatorname{\mathsf{TC}}_{m}$ of the total space of $p$ is almost thrice the value of $\operatorname{\mathsf{TC}}_{m}[\hskip 1.13809ptp:SP^{2}(M_{1})\to{\mathbb{T}}]$.
6. Symmetric products of non-orientable surfaces
Let $N_{g}$ denote the non-orientable surfaces of genus $g\geq 1$. In this section, we shall prove Theorem B, which computes the Lusternik–Schnirelmann category of the symmetric $n$-th product $SP^{n}(N_{g})$ for each $n,g\geq 1$. We will also show that the $i$-th higher homotopy group of $SP^{n}(N_{g}))$ vanishes if $n>i\geq 2$ (c.f. Theorem E).
6.A. Homotopy groups
It follows from the Dold–Thom Theorem [DT] that $SP^{\infty}(M_{g})\simeq(S^{1})^{2g}\times{\mathbb{C}}P^{\infty}$ and $SP^{\infty}(N_{g})\simeq(S^{1})^{g-1}\times{\mathbb{R}}P^{\infty}$ for each $g$. In fact, these homotopy equivalences are recovered from the following general result on homotopy splitting (see [KS, Section 6] for a proof).
6.1 Proposition.
If $X$ is a $2$-dimensional CW complex such that
$H_{1}(X;{\mathbb{Z}})={\mathbb{Z}}_{k_{1}}\oplus\cdots\oplus{\mathbb{Z}}_{k_{r}}\oplus{\mathbb{Z}}^{a}\hskip 8.53581pt\text{and}\hskip 8.53581ptH_{2}(X;{\mathbb{Z}})={\mathbb{Z}}^{b}$
for some non-negative integers $a,b,r$, and $k_{i}$, then
$SP^{\infty}(X)\simeq\left(S^{1}\right)^{a}\times\left({\mathbb{C}}P^{\infty}\right)^{b}\times L_{k_{1}}\times\cdots\times L_{k_{r}},$
where $L_{k_{i}}=S^{\infty}/{\mathbb{Z}}_{k_{i}}$ is the classifying space of the group ${\mathbb{Z}}_{k_{i}}$ for each $i$.
We use the above proposition and our discussion in Section 2.A to obtain the following result on the vanishing of certain higher homotopy groups of $SP^{n}(X)$.
6.2 Proposition.
Let $X$ be a $2$-dimensional CW complex with only one vertex, satisfying the hypothesis of Proposition 6.1. Then, for all $n\geq 3$, $\pi_{2}(SP^{n}(X))={\mathbb{Z}}^{b}$, and for each $i\geq 3$, $\pi_{i}(SP^{n}(X))=0$ for all $n>i$.
Proof.
Let $p:\widetilde{X}\to\overline{SP}^{\infty}(X)$ be the universal covering. Let us fix some $i\geq 2$. Then for any given $n\geq i+1$, if $p_{n}:Y\to\overline{SP}^{n}(X)$ is the universal covering map, then $p_{n}$ is just the pullback of $p$ with respect to the basepoint inclusion map $\overline{SP}^{n}(X)\hookrightarrow\overline{SP}^{\infty}(X)$. In view of this and (a), the $(i+1)$-skeleton of $\overline{SP}^{\infty}(X)$ coincides with the $(i+1)$-skeleton of $\overline{SP}^{n}(X)$. In particular, $\widetilde{X}^{(i+1)}=Y^{(i+1)}$. Therefore, for $n>i\geq 2$, we get
$\pi_{i}(SP^{n}(X))=\pi_{i}(\overline{SP}^{n}(X))=\pi_{i}(Y)=\pi_{i}(Y^{(i+1)})=\pi_{i}(\widetilde{X}^{(i+1)})=\pi_{i}(\widetilde{X}).$
In view of Proposition 6.1, we have that $\widetilde{X}\simeq{\mathbb{R}}^{a}\times({\mathbb{C}}P^{\infty})^{b}\times(S^{\infty})^{r}\simeq({\mathbb{C}}P^{\infty})^{b}$. The conclusion now follows directly from the fact that $\pi_{2}(({\mathbb{C}}P^{\infty})^{b})={\mathbb{Z}}^{b}$ and $\pi_{i}(({\mathbb{C}}P^{\infty})^{b})=0$ for all $i\geq 3$[Ha, Example 4.50]. ∎
We now focus on the symmetric products of surfaces and prove Theorem E.
Proof of Theorem E.
Clearly, $X=M_{g}$ satisfies the hypothesis of Proposition 6.1 with $a=2g$, $b=1$, and $r=0$. Hence, we apply Proposition 6.2 to $X=M_{g}$ and conclude that for $SP^{n}(M_{g})$, the second homotopy group is ${\mathbb{Z}}$ whenever $n\geq 3$, and $\pi_{i}(SP^{n}(M_{g}))=0$ for all $3\leq i<n$.
Similarly, $X=N_{g}$ satisfies the hypothesis of Proposition 6.1 with $r=1$, $k_{1}=2$, $a=g-1$, and $b=0$. So, we again apply Proposition 6.2 to $X=N_{g}$ and conclude that for $\pi_{i}(SP^{n}(N_{g}))=0$ for all $2\leq i<n$. ∎
The existence of spin structures on the universal covers of $SP^{n}(N_{g})$ for $n\geq 3$ can now be justified.
Proof of Corollary E.1.
We recall that for a closed $k$-manifold $M$, its universal covering admits a spin structure if and only if the evaluation of its second Stiefel–Whitney class $w_{2}(M)$ is zero on each spherical ${\mathbb{Z}}_{2}$-homology class of $M$. Since $\pi_{2}(SP^{n}(N_{g}))=0$ for $n\geq 3$ and $g\geq 1$ by Theorem E, $SP^{n}(N_{g})$ has no non-trivial spherical homology classes. Thus, the conclusion follows by the above criterion. ∎
Motivated by [DDJ, Theorem 9.12] in the case of the symmetric products of orientable surfaces, we ask the following question.
6.3 Question.
Is the orientable double cover of $SP^{n}(N_{g})$ a spin $2n$-manifold for some integers $n\geq 2$ and $g\geq 2$?
6.B. Cohomology
Let $H^{*}(X):=H^{*}(X;{\mathbb{Z}}_{2})$ and $H_{*}(X):=H_{*}(X;{\mathbb{Z}}_{2})$.
The structure of the ${\mathbb{Z}}_{2}$-cohomology ring $H^{*}(SP^{n}(N_{g}))$ in the case $n\geq 2$ was first studied by Kallel and Salvatore in [KS, Section 4.2].
The non-orientable surface $N_{g}$ can be viewed as a connected sum of $g$ copies of projective planes, say ${\mathbb{P}}_{i}$ for $1\leq i\leq g$. The CW structure on $N_{g}$ consists of one 00-cell, $g$ $1$-cells, and one $2$-cell attached by the word $e_{1}^{2}e_{2}^{2}\cdots e_{g}^{2}$. Here, $e_{i}\in H_{1}({\mathbb{P}}_{i})$ is the generator for each $i$. Using this $\Delta$-complex structure, the elements $e_{i}$ form a basis for $H_{1}(N_{g})$. We use the same notations $e_{i}$ for their images in $H_{1}(SP^{n}(N_{g}))$ under the map induced in homology by the basepoint inclusion $N_{g}\hookrightarrow SP^{n}(N_{g})$. Let $d\in H_{2}(N_{g})$ denote the ${\mathbb{Z}}_{2}$-fundamental class and its image in $H_{2}(SP^{n}(N_{g}))$. Let us denote the Hom duals of $e_{i}$ and $d$ by $e_{i}^{*}$ and $d^{*}$, respectively. Note that in $H^{*}(N_{g})$, we have that $e_{i}^{*}e_{j}^{*}=0$ for each $i\neq j$ and $(e_{i}^{*})^{2}=d^{*}$[Ha].
The following statement, which describes the ${\mathbb{Z}}_{2}$-cohomology ring $H^{*}(SP^{n}(N_{g}))$, can be seen as an analog of Theorem 5.1 in the non-orientable setting.
6.4 Theorem ([KS]).
The ${\mathbb{Z}}_{2}$-cohomology ring $H^{*}(SP^{n}(N_{g}))$ is generated by the cohomology classes $e_{1}^{*},e_{2}^{*},\ldots,e_{g}^{*}$ and $d^{*}$ under the following relations:
  1. (1)
    For each $1\leq i\leq g$, $(e_{i}^{*})^{2}=d^{*}$, and
  2. (2)
    $e_{i_{1}}^{*}\cdots e_{i_{r}}^{*}(d^{*})^{s}=0$ whenever $r+s\geq n+1$ for distinct indices $i_{1},\ldots,i_{r}$.
We note that it follows from [KS, Lemma 19] that $(d^{*})^{n}\neq 0$.
6.C. LS-category and topological complexity
We are now in a position to compute the Lusternik–Schnirelmann category of $SP^{n}(N_{g})$ for each $n$ and $g$.
Proof of Theorem B.
Of course, for each $n,g\geq 1$, we have the dimensional upper bound $\operatorname{\mathsf{cat}}(SP^{n}(N_{g}))\leq\dim(SP^{n}(N_{g}))=2n$ due to Theorem 2.1. We now obtain the lower bounds separately in the cases $n\leq g$ and $n>g$.
First, let $n\leq g$. Then we have from Theorem 6.4 that
$\left(e_{1}^{*}\right)^{2}\left(e_{2}^{*}\right)^{2}\cdots\left(e_{n}^{*}\right)^{2}=\left(d^{*}\right)^{n}\neq 0.$
Therefore, Theorem 2.1 gives the inequality $2n\leq c\ell_{{\mathbb{Z}}_{2}}(SP^{n}(N_{g}))\leq\operatorname{\mathsf{cat}}(SP^{n}(N_{g}))$ when $n\leq g$.
Next, let $n>g$. Then we again get from Theorem 6.4 that
$\left(e_{1}^{*}\right)^{2n-2g+2}\left(e_{2}^{*}\right)^{2}\left(e_{3}^{*}\right)^{2}\cdots\left(e_{g}^{*}\right)^{2}=\left(d^{*}\right)^{n-g+1}\left(d^{*}\right)^{g-1}=\left(d^{*}\right)^{n}\neq 0.$
As before, we get $2n\leq c\ell_{{\mathbb{Z}}_{2}}(SP^{n}(N_{g}))\leq\operatorname{\mathsf{cat}}(SP^{n}(N_{g}))$ when $n>g$. Hence, we have that $\operatorname{\mathsf{cat}}(SP^{n}(N_{g}))=2n$ for all $n,g\geq 1$. ∎
Unlike $\operatorname{\mathsf{dcat}}(SP^{n}(M_{g}))$, here we can’t determine $\operatorname{\mathsf{dcat}}(SP^{n}(N_{g}))$ when $n\geq 2$ because $c\ell_{{\mathbb{Z}}_{2}}(X)$ is not a lower bound to $\operatorname{\mathsf{dcat}}(X)$. The methods of [J2, Section 3] also do not help here because $\pi_{1}(SP^{n}(N_{g}))={\mathbb{Z}}^{g-1}\oplus{\mathbb{Z}}_{2}$ has non-trivial torsion.
We follow Gromov [Gr] and say that a closed $k$-manifold $M$ is essential if a map $u:M\to B\pi_{1}(M)$ that induces an isomorphism of the fundamental groups cannot be deformed to the $(k-1)$-skeleton of the CW complex $B\pi_{1}(M)$.
Proof of Corollary B.1.
Recall that a closed $k$-manifold $M$ is essential if and only if $\operatorname{\mathsf{cat}}(M)=\dim(M)=k$, see [KR, Theorem 4.1] (and also [CLOT, Theorem 2.51]). Therefore, given $n\geq 1$, $SP^{n}(M_{g})$ is essential precisely when $n\leq g$ because of Theorem 5.3, and $SP^{n}(N_{g})$ is essential for all $g\geq 1$ due to Theorem B. ∎
For the topological complexity of $SP^{n}(N_{g})$ for each $n,g\geq 1$, we have the bounds $2n\leq\operatorname{\mathsf{TC}}(SP^{n}(N_{g}))\leq 4n$ due to Theorems 2.3 and B. In the case $n=1$, the lower bound is not attained for any $g$ while the upper bound is attained for each $g\geq 2$, see [Dr2] and [CV]. We now focus on the case $n\geq 2$ and explain why the technique of Section 5.B and [DDJ, Section 6.C] that helps determine $\operatorname{\mathsf{TC}}(SP^{n}(M_{g}))$ does not help in determining $\operatorname{\mathsf{TC}}(SP^{n}(N_{g}))$ for general $n\geq 2$ and $g\geq 1$.
Given the ${\mathbb{Z}}_{2}$-cohomology classes $e_{i}^{*}\in H^{1}(SP^{n}(N_{g}))$ for each $1\leq i\leq g$, let us define the ${\mathbb{Z}}_{2}$-zero-divisors $\overline{e_{i}^{*}}=e_{i}^{*}\otimes 1+1\otimes e_{i}^{*}$. It is easy to see using Theorem 6.4 that in mod $2$, for any $k\leq g$, we have that
$\left(\overline{e_{1}^{*}}\right)^{2}\left(\overline{e_{2}^{*}}\right)^{2}\cdots\left(\overline{e_{k}^{*}}\right)^{2}=\left(\left(e_{1}^{*}\right)^{2}\otimes 1+1\otimes\left(e_{1}^{*}\right)^{2}\right)\cdots\left(\left(e_{k}^{*}\right)^{2}\otimes 1+1\otimes\left(e_{k}^{*}\right)^{2}\right)$
$=\left(d^{*}\otimes 1+1\otimes d^{*}\right)^{k}=\begin{cases}\left(d^{*}\right)^{k}\otimes 1+1\otimes\left(d^{*}\right)^{k}&\textup{if }k\textup{ is even}\\ \sum_{i=0}^{k}\left(\left(d^{*}\right)^{k-i}\otimes\left(d^{*}\right)^{i}\right)&\textup{if }k\textup{ is odd}.\end{cases}$
If $n>g$, then of course, $k\leq n$. Otherwise, if $n\leq g$, then we take $k\leq n$. In any case, because $k\leq n$, we have that $(d^{*})^{k-i}\neq 0$ and $(d^{*})^{i}\neq 0$ for all $0\leq i\leq k$ by [KS, Lemma 19]. Since we are working in the coefficient field ${\mathbb{Z}}_{2}$, we must have $(d^{*})^{k-i}=(d^{*})^{i}=1$ for each $i$. Thus,
$\left(\overline{e_{1}^{*}}\right)^{2}\left(\overline{e_{2}^{*}}\right)^{2}\cdots\left(\overline{e_{k}^{*}}\right)^{2}=\begin{cases}2&\textup{if }k\textup{ is even}\\ k+1&\textup{if }k\textup{ is odd}\end{cases}\hskip 2.84526pt=0$
in mod $2$. Hence, the ${\mathbb{Z}}_{2}$-zero-divisor cup-length of $SP^{n}(N_{g})$ for $n,g\geq 1$ does not give us a lower bound to $\operatorname{\mathsf{TC}}(SP^{n}(N_{g}))$ greater than $2n$ if we use techniques from Section 5.B. We also don’t know whether an upper bound to $\operatorname{\mathsf{TC}}(SP^{n}(N_{g}))$ smaller than $4n$ can be obtained in general. So, the topological complexity of $SP^{n}(N_{g})$ remains undetermined in the cases $n\geq 2$. Note that in the case $n=1$, $\operatorname{\mathsf{TC}}_{m}(N_{g})$ is completely determined for all $m\geq 2$ and $g\geq 1$[GGGL][Dr2][CV]. Concretely, $\operatorname{\mathsf{TC}}_{m}(N_{g})=2m$ for each $m$ and $g$, except for the case $(m,g)=(2,1)$ when $\operatorname{\mathsf{TC}}(N_{1})=3$.
6.5 Remark.
Since $SP^{n}(N_{1})=SP^{n}({\mathbb{R}}P^{2})={\mathbb{R}}P^{2n}$ for each $n\geq 1$, any general result on the topological complexity of $SP^{n}(N_{g})$ for $n\geq 2$ must subsume the main result of [FTY], which says that $\operatorname{\mathsf{TC}}({\mathbb{R}}P^{2n})$ is equal to the immersion dimension of ${\mathbb{R}}P^{2n}$. This indicates that determining $\operatorname{\mathsf{TC}}(SP^{n}(N_{g}))$ can be very difficult.
Acknowledgement
The author would like to thank Alexander Dranishnikov for suggesting the problem of determining the sequential topological complexities of symmetric products of orientable surfaces.